PREPARATION OF PAPERS FOR JACoW CONFERENCES^{∗}

J. Poole, C. Petit-Jean-Genaz, CERN, Geneva, Switzerland

C. Eyberger^{†}, ANL, Argonne, IL 60439, USA

V. R.W. Schaa, GSI, Darmstadt, Germany T. Satogata, JLab, Newport News, VA 23606, USA

I. Andrian, Elettra, Trieste, Italy

Abstract

Many conference series have adopted the same standards for electronic publication and have joined the Joint Accelerator Conferences Website (JACoW) collaboration for the publication of their proceedings. This document describes the common requirements for the submission of papers to these conferences. Please consult individual conference information for page limits, method of electronic submission, etc. It is not intended that this should be a tutorial in word processing; the aim is to explain the particular requirements for electronic publication at these conference series.

SUBMISSION OF PAPERS

Each author should submit the PDF file and all of the source files (text and figures), to enable the paper to be reconstructed if there are processing difficulties.

MANUSCRIPTS

Templates are provided for recommended software and authors are advised to use them. Please consult the individual conference help pages if questions arise.

General Layout

These instructions are a typical implementation of the requirements. Manuscripts should have:

• Either A4 (21.0cm × 29.7cm; 8.27in × 11.69in) or US letter size (21.6cm × 27.9cm; 8.5in × 11.0in) paper.

• Single-spaced text in two columns of 82.5mm (3.25in) with 5.3mm (0.2in) separation. More recent versions of MSWord have a default spacing of 1.5 lines; authors must change this to 1 line.

• The text located within the margins specified in Table 1 to facilitate electronic processing of the PDF file.

The layout of the text on the page is illustrated in Fig. 1. Note that the paper’s title and the author list should be the width of the full page. Tables and figures may span the whole 170mmpage width, if desired (see Fig. 2), but if they span both columns, they should be placed at either the top or bottom of a page to ensure proper flow of the text (Word templates only: the text should flow from top to bottom in each column).

Table 1: Margin Specifications

Margin  A4 Paper  US Letter Paper
Top  37mm (1.46in) 0.75in (19mm)
Bottom 19mm (0.75in) 0.75in (19mm)
Left  20mm (0.79in) 0.79in (20mm)
Right  20mm (0.79in) 1.02in (26mm)

Figure 1: Layout of papers.

Fonts

In order to produce good Adobe Acrobat PDF files, authors using a L^{A}TEX template are asked to use only Times (in roman [standard], bold or italic) and symbols from the standard set of fonts. In Word use only Symbol and, depending on your platform, Times or Times New Roman fonts in standard, bold or italic form.

Title and Author List

The title should use 14pt bold uppercase letters and be centered on the page. Individual letters may be lowercase to avoid misinterpretation (e.g., mW, MW). To include a funding support statement, put an asterisk after the title and the support text at the bottom of the first column on page 1— in Word, use a text box; in L^{A}TEX, use \thanks.

The names of authors, their organizations/affiliations and mailing addresses should be grouped by affiliation and listed in 12pt upper- and lowercase letters. The name of the submitting or primary author should be first, followed by the co-authors, alphabetically by affiliation.

Section Headings

Section headings should not be numbered. They should use 12pt bold uppercase letters and be centered in the column. All section headings should appear directly above the text—there should never be a column break between a heading and the following paragraph.

Subsection Headings

Subsection headings should not be numbered. They should use 12pt italic letters and be left aligned in the column. Subsection headings use Title Case (or Initial Caps) and should appear directly above the text—there should never be a column break between a heading and the following paragraph.

∗ Work supported by ...

† cee@aps.anl.gov

Figure 2: Example of a full-width figure showing the JACoW Team at their annual meeting in 2012. This figure has a multi-line caption that has to be justified rather than centered.

Third-level Headings are created with the L^{A}TEX command \subsubsection. In the Word templates authors must bold the text themselves; this heading should be used sparingly. See Table 2 for its style details.

Paragraph Text

Paragraphs should use 10pt font and be justified (touch each side) in the column. The beginning of each paragraph should be indented approximately 3mm (0.13in). The last line of a paragraph should not be printed by itself at the beginning of a column nor should the first line of a paragraph be printed by itself at the end of a column.

Figures, Tables and Equations

Place figures and tables as close to their place of mention as possible. Lettering in figures and tables should be large enough to reproduce clearly. Use of non-approved fonts in figures can lead to problems when the files are processed. L^{A}TEX users should be sure to use non-bitmapped versions of Computer Modern fonts in equations (Type 1 PostScript or OpenType fonts are required, Their use is described in the JACoW help pages [1]).

Each figure and table must be numbered in ascending order (1, 2, 3, etc.) throughout the paper. Figure captions (10pt font) are placed below figures, and table captions are placed above tables. Single-line captions are centered in the column, while captions that span more than one line should be justified. The L^{A}TEX template uses the ‘booktabs’ package to format tables.

A simple way to introduce figures into a Word document is to place them inside a table that has no borders. This is done in Word as described below.

Note: If the figure spans both columns, do all steps. If the figure is contained in a single column, start at step 5.

1. Insert a continuous section break.

2. Insert two empty lines (makes later editing easier).

3. Insert another continuous section break.

4. Click between the two section breaks and Page Layout → Columns → Single.

5. Insert → Table select a one-column, two-row table.

6. Paste the figure in the first row of the table and adjust the size as appropriate.

7. Paste/Type the caption in the second row and apply the appropriate figure caption style.

8. Table → Table properties → Borders and Shading → None.

9. Table → Table properties → Alignment → Center.

10. Table → Table properties → Text wrapping → None.

11. Remove blank lines from in and around the table.

12. If necessary play with the cell spacing and other parameters to improve appearance.

If a displayed equation needs a number (i. e., it will be referenced), place it flush with the right margin of the column (see Eq. ??). The equation itself should be indented (centered, if possible). Units should be written using the roman (standard) font, not the italic font:

C_{B} = \frac{q^{3}}{3ϵ_{0}mc} = 3.54µeV/T  (1)

References

All bibliographical and web references should be numbered and listed at the end of the paper in a section called REFERENCES. When citing a reference in the text, place the corresponding reference number in square brackets [2]. The reference citations in the text should be numbered in ascending order.

A URL may be included as part of a reference, but its hyperlink should NOT be added. Multiple citations should appear in the same square bracket [1, 3, 4] or with ranges, e.g., [1]– [4] or [1–4, 10].

Examples of correctly formatted references can be found at the JACoW website (http://JACoW.org). Once there, click on the ‘for Authors’ link at the top and then on the ‘Formatting Citations’ link in the left-hand column [1].

Footnotes

Footnotes on the title and author lines may be used for acknowledgments, affiliations and e-mail addresses. A nonnumeric sequence of characters (*, #, †, ‡) should be used. Word users—DO NOT use Word’s footnote feature (Insert, Footnote) to insert footnotes, as this will create formatting problems. Instead, insert the title or author footnotes manually in a text box at the bottom of the first column with a line at the top of the text box to separate the footnotes from the rest of the paper’s text. The easiest way to do this is to copy the text box from the JACoW template and paste it into your own document. These “pseudo footnotes” in the text box should only appear at the bottom of the first column on the first page.

Any other footnote in the body of the paper should use the normal numeric sequencing and appear as footnote^{1} in the same column where they are used.

Acronyms

Acronyms should be defined the first time they appear.

1 This text should appear in the column where it was referenced.

STYLES

Table 2 summarizes the fonts and spacings used in the styles of a JACoW template (these are implemented in the L^{A}TEX class file).

Table 2: Summary of Styles

Style  Font  Space Space
Before After
14pt  0pt 3pt
Paper Title Upper case except for

required lower case letters

Bold

Author list 12pt  9pt 12pt

Upper and Lower case

Section  12pt  9pt 3pt
Heading Uppercase

bold

Subsection 12pt  6pt 3pt
Heading Initial Caps

Italic

Third-level 10pt  6pt 0pt
Heading Initial Caps

Bold

Figure  10pt  3pt 6pt

Captions

Table  10pt  3pt 3pt

Captions

Equations 10pt base font  12pt 12pt
References 9/10pt, justified with 0.25in
hanging indent, reference  ≥0pt ≥0pt

number right aligned

PAGE NUMBERS

DO NOT include any page numbers. They will be added when the final proceedings are produced.

TEMPLATES

Templates and examples can be retrieved through web browsers such as Firefox, Chrome and Internet Explorer by saving to disk.

Template documents for the recommended word processing software are available from the JACoW website (http: //JACoW.org) and exist for L^{A}TEX, Microsoft Word (Mac and PC) and OpenOffice for US letter and A4 paper sizes.

Use the correct template for your paper size and version of Word. Do not transport Microsoft Word documents across platforms, e.g., Mac ↔ PC. When saving a Word 2010 file (PC), be sure to click ‘Embed fonts’ in the Save options. Fonts are embedded by default when printing to PDF on Mac OSX.

Please see the information and help files for authors on the JACoW.org website for instructions on how to install templates in your Microsoft templates folder.

CHECKLIST FOR ELECTRONIC PUBLICATION

• Use only Times or Times New Roman (standard, bold or italic) and Symbol fonts for text—10pt minimum except references, which can be 9pt or 10pt.

• Figures should use Times or Times New Roman (standard, bold or italic) and Symbol fonts when possible— 6pt minimum.

• Check that citations to references appear in sequential order and that all references are cited [10].

• Check that the PDF file prints correctly.

• Check that there are no page numbers.

• Check that the margins on the printed version are within ±1 mm of the specifications.

• L^{A}TEX users can check their margins by invoking the boxit option.

CONCLUSION

Any conclusions should be in a separate section directly preceding the ACKNOWLEDGMENT, APPENDIX, or REFERENCES sections, in that order.

ACKNOWLEDGMENT

Any acknowledgment should be in a separate section directly preceding the REFERENCES or APPENDIX section.

APPENDIX

Any appendix should be in a separate section directly preceding the REFERENCES section. If there is no REFER-ENCES section, this should be the last section of the paper.

REFERENCES

[1] JACoW.org website: http://jacow.org/index.php?n=

Authors.UsingLaTeX {no hyperlink, no period after URL}

[2] A.N. Other, “A Very Interesting Paper”, EPAC’96, Sitges,

June 1996, MOPCH31 (1996), http://www.JACoW.org {no hyperlink, no period after URL}

[3] F.E. Black et al., This is a Very Interesting Book, (New York: Knopf, 2007), 52.

[4] G.B. Smith et al., “Title of Paper,” IPAC’14, Dresden, Germany, June 2014, MOXAP07, These Proceedings.

⋮

[10] B. Gnats, A. Jones, Phys. Rev. ST Accel. Beams 1, 011502